OpenStudy (anonymous):

find the polynomial f(x) that has the roots -2, 3 multiplicity of 2

OpenStudy (anonymous):

Roots essentially mean where x intersects the x-axis or where y=0 x=-2 x=3 multiplicity 2 for x=-2, you can do (x+2) = f(x) for x = 3, you can do (x-3) = f(x) for multiplicities, you simply have to raise the power for x=3 multiplicity 2, you do (x-3)^2 = f(x) Combine these to get your equation: \[f(x) = (x+2)(x-3)^2\] multiply and simplify to get your polynomial

OpenStudy (anonymous):

Thank you so much!